OpenStudy (anonymous):

MEDAL <3 A system of equations is shown below: 3x – y = 2 x + y = 6 The x-coordinate of the solution to this system of equations is

OpenStudy (helder_edwin):

add both equations

OpenStudy (helder_edwin):

what do u get?

OpenStudy (anonymous):

3x = 8

OpenStudy (anonymous):

2.67 ?

OpenStudy (the_fizicx99):

3x + x = ? Remember that x + x + x = 3x x * x * x = x^3 There's a difference :)

OpenStudy (anonymous):

So 4x? @tHe_FiZiCx99

OpenStudy (the_fizicx99):

 Yes, 4x. The -y + y cancel out.